OpenStudy (anonymous):

Consider a triangle where a=3, b=5 and c=7. Use the law of cosines to find angle C. possible answers are 60 deg, 120 deg, 135 deg or not possible. I came up with 3.97 so I must have did this wrong. Pls help

OpenStudy (anonymous):

As long as you are consistent you will be fine so \[c ^{2}=a^{2}+b^{2}-2abCos(C)\] \[7^{2}=3^{2}+5^{2}-2*3*5*Cos(C)\] \[49=9+25-30*Cos(C)\] \[15=-30*Cos(C)\] \[-1/2=Cos(C)\] \[Cos^{-1}(-1/2)=C\] So, C=120.

OpenStudy (anonymous):

Oh, thank you so much! I was lost on this. For some reason your response didn't show up in notifications and i just clicked on my question and your response was there. Am I suppose to refresh my screen often?